OpenStudy (anonymous):

What is the slope of the line determined by any two solutions to the equation 2/x+3/y=0 ? Express your answer as a common fraction.

OpenStudy (anonymous):

\[\Large \begin{array}{l} \frac{2}{x} + \frac{3}{y} = 0\\ \frac{3}{y} = - \frac{2}{x}\\ y = \frac{{ - 3x}}{2} \end{array}\]